# black holes; early universe; general relativity

### v1

@@ Abstract
Primordial black holes (PBHs) may have formed through the gravitational collapse of cosmological perturbations that were generated and stretched during the inflationary era, later entering the cosmological horizon during the decelerating phase, if their amplitudes were sufficiently large. In this article, I will briefly introduce the basic concept of PBHs and review the formation dynamics through this mechanism, the estimation of the initial spins of PBHs and the time evolution of type II fluctuations, with a focus on the radiation-dominated and (early) matter-dominated phases.
@@ 1 Introduction
Black holes may have formed in the early Universe. This possibility was first considered by Zel’dovich and Novikov (1966) Zel’dovich and Novikov (1967) and Hawking (1971) Hawking (1971). These black holes have been termed primordial black holes (PBHs). The observational relevance of PBHs to cosmology has been established by the subsequent work of Carr and Hawking (1974) Carr and Hawking (1974) and Carr (1975) Carr (1975).
@@
The motivation for PBH research is multi-faceted. Since PBHs can, in principle, be observed, they provide valuable information about the early Universe. In this sense, PBHs can be regarded as the fossils of the early Universe. Furthermore, they are also one of the most promising candidates for dark matter. Moreover, Hawking evaporation plays an important role in the study of PBHs. Hawking recognised PBHs as a unique laboratory for studying black hole evaporation when he discovered this phenomenon in 1974 Hawking (1974, 1975). Hawking evaporation is key to understanding quantum gravity, where the information loss problem has been extensively discussed. The details of this evaporation may depend on quantum gravity and high-energy physics. PBHs are also important candidates for gravitational wave sources. Recently, several gravitational wave observatories have become available as observational tools to study strongly gravitating but otherwise dark objects, following LIGO’s first direct detection of gravitational waves in 2015 Abbott et al. (2016). It has been realised that gravitational waves provide a unique means of probing PBHs. This suggests that PBHs lie at the intersection of various growing fields of modern physics, such as cosmology, general relativity, gravitational waves, quantum gravity, and high-energy physics.
@@
As observational data accumulate, gravitational waves are becoming an increasingly important means of observing our Universe. More than 100 events have been observed by the LIGO-Virgo-KAGRA (LVK) collaboration, and many of these have been identified as binary black holes. Several groups proposed the possibility that these binary black holes may be of cosmological origin, as many of them are approximately 30 $M_{\odot}$, which is more massive than was typically expected based on the standard theory of stellar evolution before the direct observations Sasaki et al. (2016); Bird et al. (2016); Clesse and García-Bellido (2017). As more observational data accumulate, information about not only the masses but also the spins of binary black holes is obtained for some events Abbott et al. (2017), which may provide a new clue to the origin of these binary black holes. From the obtained mass function of binary black holes, a search for the population of PBHs was carried out Franciolini et al. (2022). It has also been argued that the existence of subsolar candidates in LIGO data is a smoking gun for PBHs because no astrophysical scenario has yet been established to form subsolar-mass black holes Carr et al. (2024). Most recently, the NANOGrav collaboration reported evidence for nanohertz gravitational waves Agazie et al. (2023), which might be consistent with gravitational waves induced by scalar perturbations that could have produced the amount of PBHs responsible for the binary black holes observed by the LVK collaboration (e.g. Afzal et al. (2023)).
@@
I would also comment on the possibility of PBHs constituting all or a considerable fraction of dark matter. We usually discuss this in terms of the fraction of PBHs to cold dark matter (CDM), $f(M)=\Omega_{\rm PBH}(M)/\Omega_{\rm CDM}$, as a function of the mass of PBHs, $M$. This fraction is observationally constrained very severely in some mass ranges but not in others. See Fig. 10 of Carr, Kohri, Sendouda and Yokoyama (2021) Carr et al. (2021) for an overview of the constraints. Recent observational constraints indicate two intriguing windows for dark matter. One is $M\sim 10^{17}-10^{23}$ g, where $f(M)$ is virtually unconstrained. This implies that PBHs could account for all the CDM for this mass range. The other mass range is $M\sim 1-10^{3}M_{\odot}$, where $f(M)\lesssim 0.1$, which is of interest in the context of terrestrial gravitational wave observations. In the latter range, however, a more recent paper has placed a new bound of $f\lesssim 10^{-3}$ for $M\sim 1-200M_{\odot}$ from LVK O3 data  Andrés-Carcasona et al. (2024).
Although there is currently a mass window in which all the CDM might be explained by PBHs, a stricter constraint could be placed on this mass window in the near future. Even if this turns out to be the case, it does not imply that studies of PBHs are without value. In this context, I would like to quote a profound statement by Bernard Carr: “Indeed their study may place interesting constraints on the physics relevant to these areas even if they never formed” Carr (2003).
@@ 2.1 Mass
The striking feature of PBHs is that they can have a large range of possible mass scales from $10^{-5}$ g to $10^{56}$ g depending on the formation scenarios at least in principle. Although the mass of the PBH may depend on the scenario, we usually assume that it can be approximately given by the mass enclosed within the cosmological horizon or Hubble horizon at the formation time $t_{f}$ from the big bang as
@@
Table 1: The cosmological time and the initial mass of PBHs if they are formed then
@@
It is nontrivial that how the initial mass of PBHs is related to the mass during their evolution. As for the mass accretion, earlier works suggest that it does not so significantly affect the mass at least in radiation domination Carr and Hawking (1974). This was confirmed by numerical simulations (e.g. Escrivà (2020)). On the other hand, the Hawking evaporation is considered to decrease the mass of PBHs. The mass loss is considered to virtually be negligible if the mass of the PBH is much more massive than the critical mass $\sim 10^{15}$ g, which will be discussed below.
@@ 2.2 Evaporation
where $\hbar$ and $k$ are the reduced Planck constant and the Boltzmann constant, respectively. This is called the Hawking evaporation. If we assume that the black hole loses its mass due to this radiation of quantum fields, which is the so-called semi-classical approximation, we obtain
@@
From Eq. (5), the critical mass is approximately given by $10^{15}$ g for which $t_{\rm ev}=t_{0}$ is satisfied, where $t_{0}$ is the age of the Universe. So, if $M\lesssim 10^{15}$ g, PBHs have dried up after the explosion until now, whether they leave Planck mass relics or not. If $M\sim 10^{15}$ g, PBHs are currently emitting X rays and $\gamma$ rays and can be observed through those emissions. If $M\gtrsim 10^{17}$ g, the evaporation is mostly negligible and the mass of the PBH remains almost constant until now.
@@ 2.3 Probability
To discuss the formation, the fraction $\beta(M)$ of the Universe which goes into PBHs when the mass contained within the cosmological horizon is $M$, is often used. This can also be regarded as the formation probability of PBHs. If we consider PBHs formed in the radiation-dominated phase, PBHs act as nonrelativistic particles surrounded by relativistic particles. Therefore, the energy density of PBHs, $\rho_{\rm PBH}$, decays as $a^{-3}$, while the energy density of relativistic particles, $\rho_{\rm rad}$, decays as $a^{-4}$, where $a$ is the scale factor of the Universe. As a result, $\beta(M)$ is not equal to the current fraction of PBHs of mass $M$ to all the CDM, which is given by
$f(M)=\left.\frac{\Omega_{\rm PBH}}{\Omega_{\rm CDM}}\right|_{t=t_{0}}.$ (6)
Taking in account the effect that PBHs are concentrated during the radiation-dominated era, we find
@@
for $M\gtrsim 10^{15}$ g. Thus, only a tiny probability $\beta\simeq 2\times 10^{-17}$ is enough to explain all of the dark matter if $M=10^{17}$ g.
@@ 3.1 Overview
The serious study of PBH formation date back to Carr (1975) Carr (1975). One of its aims is to theoretically predict $\beta(M)$ and therefore $f(M)$ and other observables of PBHs for a given cosmological scenario. We can obtain the information of the early Universe from observational data on PBHs only through these studies. Apart from such an observational motivation, through the study of formation we can understand physics in PBH formation and investigate new phenomena and/or new physics in highly nonlinear general relativistic dynamics and high-energy physics. There have been a lot of scenarios proposed for PBH formation. One of the most standard ones is the direct collapse of a large amplitude of primordial perturbations generated by inflation. This can be considered as ‘inevitable’ as the inflationary cosmology has been regarded as an essential part of standard cosmology. Among the other alternative scenarios, domain wall collapse, bubble nucleation, collapse of string networks and phase transitions have attracted more attention than others. Hereafter, we will focus on the formation scenario from fluctuations generated by inflation. Studies on this scenario have developed more than on other scenarios. The key ideas to this scenario is the following: fluctuations generated by inflation, long-wavelength solutions, formation threshold, black hole critical behaviour, dependence on the equation of state (EOS) and statistics on the abundance estimate. We will briefly view these topics below.
@@ 3.2 Fluctuations generated by inflation
The most striking feature of inflation is that it can not only solve the flatness problem and the horizon problem but also provide the mechanism to generate fluctuations as quantum effects. Those fluctuations seed structure formation of different scales in our Universe and provide anisotropy in cosmic microwave background (CMB) currently observed with high accuracy.
Figure 1: Evolution of the scale of fluctuations after generated by inflation
The evolution of fluctuations in the inflationary cosmology is schematically illustrated in Fig. 1. The quantum fluctuations generated by inflation are stretched to the length scales much larger than the Hubble horizon scale $cH^{-1}$ because the fluctuation scale, which is proportional to $a$, expands faster than the Hubble horizon scale, which is proportional to $t$, due to the accelerated expansion. This process is also considered to lose quantum coherence from the fluctuations. The generated classical fluctuations are further stretched away compared to the Hubble length as long as the inflationary phase continues. After the inflation ends, the expansion of the Universe begins to be decelerated. There is a possibility that the Universe might experience the early matter-dominated phase due to the harmonic oscillation of an inflaton field. In any case, the Universe eventually gets dominated by the radiation field, the tightly coupled relativistic particles in an almost complete thermal equilibrium state. This process is called reheating. If we consider the scale of the fluctuation much larger than the Hubble horizon in the decelerated Universe, the scale of the fluctuation, which is proportional to $a$, expands slower than the Hubble horizon scale, which is proportional to $t$. The time when the fluctuation scale gets as large as the Hubble scale is called the horizon entry of the perturbation. Although the inflationary cosmology has been becoming an essential part of the standard cosmology, there has been no standard inflation model until now. There are lots of inflation models, each of which gives the power spectrum $P_{\zeta}(k)$ and the other statistics of the curvature perturbations $\zeta$. See Ref. Sasaki et al. (2018) for details of different inflation models in the context of PBH formation.
@@ 3.3 Large-amplitude long-wavelength solutions
As discussed above, the scale of fluctuation generated by inflation gets much larger than the Hubble horizon scale. Therefore, after inflation ends, we have the fluctuations of super-horizon scales in the decelerated expansion. In this phase, we can construct the so-called long-wavelength solutions of the Einstein equation Tomita (1972); Shibata and Sasaki (1999); Lyth et al. (2005); Polnarev and Musco (2007).
Figure 2: Schematic figure of long-wavelength solutions
In order to discuss the PBH formation in radiation domination, it is necessary to deal with a nonlinearly large amplitude of perturbation. We implement gradient expansion of the Einstein equation to obtain the long-wavelength solutions, where the spatial derivative is regarded as much smaller than the time derivative. More precisely, we take the following procedure Shibata and Sasaki (1999); Harada et al. (2015). We first decompose the metric to the following $3+1$ form:
$ds^{2}=-\alpha^{2}dt^{2}+e^{2\zeta}a^{2}(t)\tilde{\gamma}_{ij}(dx^{i}+\beta^{i}dt)(dx^{j}+\beta^{j}dt),$ (8)
where $i,j$ run over 1,2 and 3 and $\alpha$, $\beta^{i}$, $\tilde{\gamma}_{ij}$ and $\zeta$ are the functions of $t$ and $x^{i}$ and called the lapse function, the shift vector, the spatial metric and the curvature perturbation, respectively, while $a(t)$ is the scale factor of the reference Friedmann-Lemaître-Robertson-Walker (FLRW) spacetime. We require that $\det(\tilde{\gamma}_{ij})=\det(\eta_{ij})$, where $\eta_{ij}$ is the time-independent metric of the flat 3 space. Then, we assume that the scale of the perturbation is much larger than the Hubble horizon scale, i.e., that
@@
is much smaller than unity. Applying the gradient expansion for the Einstein equation, we can obtain growing-mode solutions with an assumption that the zeroth order solutions in powers of $\epsilon$ takes the following form:
$ds^{2}=-dt^{2}+e^{2\zeta_{0}({\bf x})}a^{2}(t)\eta_{ij}dx^{i}dx^{j}.$ (10)
Note that this assumption is compatible with the comoving slice, the constant-mean-curvature slice and the uniform-density slice but not with the conformal Newtonian gauge condition, which is often used in cosmology. The higher-order terms of the solutions are obtained in terms of the function $\zeta_{0}({\bf x})$ if we impose the appropriate gauge conditions. In other words, $\zeta_{0}({\bf x})$, which is not suppressed due to the long-wavelength scheme, generates the long-wavelength solutions. For example, the density perturbation $\delta$ is $O(\epsilon^{2})$ and is given by
@@
$ds^{2}=-dt^{2}+e^{2\zeta_{0}(r)}[dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})],$ (12)
in the conformally flat coordinates, we can show that this is equivalent to the asymptotically quasi-homogeneous solutions developed in Ref. Polnarev and Musco (2007) in the Misner-Sharp formulation, where the comoving slice and the comoving thread are adopted, and the explicit transformation between them is given in Ref. Harada et al. (2015).

### v3

@@ Abstract
Primordial black holes (PBHs) may have formed through the gravitational collapse of cosmological perturbations that were generated and stretched during the inflationary era, later entering the cosmological horizon during the decelerating phase, if their amplitudes were sufficiently large. In this review paper, we will briefly introduce the basic concept of PBHs and review the formation dynamics through this mechanism, the estimation of the initial spins of PBHs and the time evolution of type II fluctuations, with a focus on the radiation-dominated and (early) matter-dominated phases.
@@ 1 Introduction
Black holes may have formed in the early Universe. This possibility was first considered by Zel’dovich and Novikov (1966) Zel’dovich and Novikov (1967) and Hawking (1971) Hawking (1971). These black holes have been termed primordial black holes (PBHs). The observational relevance of PBHs to cosmology has been established by the subsequent work of Carr and Hawking (1974) Carr and Hawking (1974) and Carr (1975) Carr (1975). See Refs. Carr and Green (2024); Riotto and Silk (2024) for the recent brief reviews of the history and the future of PBHs.
Although our Universe can be well approximated by a homogeneous and isotropic cosmological model – the Friedmann-Lemaître-Robertson-Walker (FLRW) solution – on large scales, cosmological structures such as galaxy clusters, galaxies, planetary systems, and stars must still form. These structures arise from primordial fluctuations. Observations of anisotropies in the cosmic microwave background (CMB) radiation have revealed that the amplitude of these primordial fluctuations is generally very small. One of the most promising mechanisms to generate such fluctuations is inflation. In this scenario, quantum fluctuations provide the origin of the perturbations, strongly suggesting that the probability of generating large-amplitude perturbations would be small but finite. During inflation, these perturbations expand along with the inflationary expansion of the Universe. After inflation, when the Universe is reheated, sufficiently large-amplitude perturbations could eventually overcome cosmic expansion and collapse into black holes due to their self-gravity. Unlike the formation of typical astrophysical black holes, the formation of PBHs does not require stellar evolution.
The formation of PBHs through primordial fluctuations was initially proposed by Hawking (1971) Hawking (1971) and Carr and Hawking (1974) Carr and Hawking (1974) and further developed by Carr (1975) Carr (1975). This formation mechanism has since been confirmed through numerical simulations based on numerical relativity, pioneered by Nadezhin, Novikov and Polnarev (1978) Nadezhin et al. (1978). In this review paper, we will explore the recent developments in PBH formation studies and their implications for cosmology.
@@
The motivation for PBH research is multi-faceted. Since PBHs can, in principle, be observed, they provide valuable information about the early Universe. In this sense, PBHs can be regarded as the fossils of the early Universe. Furthermore, they are also one of the most promising candidates for dark matter. Moreover, Hawking evaporation plays an important role in the study of PBHs. Hawking recognised PBHs as a unique laboratory for studying black hole evaporation when he discovered this phenomenon in 1974 Hawking (1974, 1975). Hawking evaporation is key to understanding quantum gravity, where the information loss problem has been extensively discussed. The details of this evaporation may depend on quantum gravity and high-energy physics. Recently, several gravitational wave observatories have become available as observational tools to study strongly gravitating but otherwise dark objects, following LIGO’s first direct detection of gravitational waves in 2015 Abbott et al. (2016). It has been realised that gravitational waves provide a unique means of probing PBHs Sasaki et al. (2018). This suggests that PBHs lie at the intersection of various growing fields of modern physics, such as cosmology, general relativity, gravitational waves, quantum gravity and high-energy physics.
@@
As observational data accumulate, gravitational waves are becoming an increasingly important means of observing our Universe. More than 100 events have been observed by the LIGO-Virgo-KAGRA (LVK) collaboration, and many of these have been identified as binary black holes. Several groups proposed the possibility that these binary black holes may be of cosmological origin, as many of them are approximately 30 $M_{\odot}$, which is more massive than was typically expected based on the standard theory of stellar evolution before the direct observations Sasaki et al. (2016); Bird et al. (2016); Clesse and García-Bellido (2017). As more observational data accumulate, information about not only the masses but also the spins of binary black holes is obtained for some events Abbott et al. (2017), which may provide a new clue to the origin of these binary black holes. From the obtained mass function of binary black holes, a search for the population of PBHs was carried out Franciolini et al. (2022). It has also been argued that the existence of subsolar candidates in LIGO data would be a smoking gun for PBHs because no astrophysical scenario has yet been established to form subsolar-mass black holes Carr et al. (2024). Most recently, the NANOGrav collaboration reported evidence for nanohertz gravitational waves Agazie et al. (2023), which might be consistent with gravitational waves induced by scalar perturbations that could have produced the amount of PBHs responsible for the binary black holes observed by the LVK collaboration (e.g. Afzal et al. (2023)). The recent observations with the James Web Space Telescope have identified several very massive galaxies at high redshifts, which might indicate a tension with predictions in the standard $\Lambda$CDM model but can be explained if we include PBH clusters in the formation scenario Liu and Bromm (2022); Hütsi et al. (2023); Yuan et al. (2024).
@@
I would also comment on the possibility of PBHs constituting all or a considerable fraction of dark matter. We usually discuss this in terms of the fraction of PBHs to cold dark matter (CDM), $f(M)=\Omega_{\rm PBH}(M)/\Omega_{\rm CDM}$, as a function of the mass of PBHs, $M$. This fraction is observationally constrained very severely in some mass ranges but not in others. See Fig. 10 of Carr, Kohri, Sendouda and Yokoyama (2021) Carr et al. (2021) for an overview of the constraints. Recent observational constraints indicate two intriguing windows for dark matter. One is $M\sim 10^{17}-10^{23}$ g, where $f(M)$ is virtually unconstrained. This implies that PBHs could account for all the CDM for this mass range. The other mass range is $M\sim 1-10^{3}M_{\odot}$, where $f(M)\lesssim 0.1$, which is of interest in the context of terrestrial gravitational wave observations. Although there is currently a mass window in which all the CDM might be explained by PBHs, a stricter constraint could be placed on this mass window in the near future. Even if this turns out to be the case, it does not imply that studies of PBHs are without value. In this context, I would like to quote a profound statement by Bernard Carr: “Indeed their study may place interesting constraints on the physics relevant to these areas even if they never formed” Carr (2003).
The aim of this article is to briefly review theoretical studies of PBH formation from primordial fluctuations, the recent estimation of the spins of PBHs and the formation of PBHs generated from the so-called type II configurations. In Sec. 2, we present the basic concept of PBHs. We discuss their masses, the Hawking evaporation, the fraction of their contribution to the whole dark matter and the formation probability. In Sec. 3, we discuss the formation of PBHs. We focus on the formation mechanism from fluctuations generated by inflation and present key ideas on this scenario. In Sec 4, we discuss the spins of PBHs. The spins of PBHs have often been regarded just as negligible until recently. We quantitatively discuss the spins of PBHs just formed not only in a radiation-dominated era but also in an (early) matter-dominated era. In Sec. 5, we discuss the initial configurations featured with their spatial geometry with throat structure, which is called type II, and the formation of PBHs from them based on the result of the recently conducted numerical simulations. In Sec. 6, we conclude the paper.
@@ 2.1 Mass
The striking feature of PBHs is that they can have a large range of possible mass scales from $\sim 10^{-5}$ g to $\sim 10^{56}$ g depending on the formation scenarios at least in principle. Although the mass of the PBH may depend on the scenario, we usually assume that it can be approximately given by the mass enclosed within the cosmological horizon or Hubble horizon at the formation time $t_{f}$ from the big bang as
@@
Table 1: The cosmological time from big bang and the initial mass of PBHs if they are formed then.
@@
The mass of PBHs may have changed in time after they formed. As for the mass accretion, earlier works suggest that it does not so significantly affect the mass at least in radiation domination Carr and Hawking (1974). This was confirmed by numerical simulations (e.g. Escrivà (2020)). On the other hand, the Hawking evaporation is considered to decrease the mass of PBHs. The mass loss is considered to be virtually negligible if the mass of the PBH is much more massive than the critical mass $\sim 10^{15}$ g, which will be discussed below.
@@ 2.2 Evaporation
where $\hbar$ and $k$ are the reduced Planck constant and the Boltzmann constant, respectively. This is called the Hawking evaporation. If we assume that the black hole loses its mass due to this radiation of quantum fields according to the Stefan-Boltzmann law, which is the so-called semi-classical approximation, we obtain
@@
From Eq. (5), the critical mass is approximately given by $10^{15}$ g for which $t_{\rm ev}=t_{0}$ is satisfied, where $t_{0}$ is the age of the Universe. So, if $M\lesssim 10^{15}$ g, PBHs have dried up after the explosion until now, whether they leave Planck mass relics or not. If $M\sim 10^{15}-10^{17}$ g, PBHs are currently emitting X rays and $\gamma$ rays and can be observed through those emissions. If $M\gtrsim 10^{17}$ g, the evaporation is mostly negligible and the mass of the PBH remains almost constant until now. It should also be noted that very recently there has been a debate on a specific hypothetical scenario that assumes that the Hawking evaporation be strongly affected due to the back reaction effect even if the black hole is macroscopic Dvali et al. (2020).
@@ 2.3 Probability
To discuss the formation, the fraction $\beta(M)$ of the Universe which goes into PBHs when the mass contained within the cosmological horizon is $M$, is often used. This can also be regarded as the formation probability of PBHs. If we consider PBHs formed in the radiation-dominated phase, PBHs act as nonrelativistic particles surrounded by relativistic particles. Therefore, the energy density of PBHs, $\rho_{\rm PBH}$, decays as $a^{-3}$, while the energy density of relativistic particles, $\rho_{\rm rad}$, decays as $a^{-4}$, where $a$ is the scale factor of the Universe. This implies $\rho_{\rm PBH}/\rho_{\rm rad}\propto a$, that is, PBHs are condensed during the radiation-dominated era in proportion to the scale factor. Taking this concentration effect into account, $\beta(M)$ is related to the the current fraction of PBHs of mass $M$ to all the CDM
$f(M)=\left.\frac{\Omega_{\rm PBH}}{\Omega_{\rm CDM}}\right|_{t=t_{0}}$ (6)
through
@@
for $M\gtrsim 10^{15}$ g Carr (1975); Carr et al. (2021). This formula can derived by the expansion law $a(t)\propto t^{1/2}$ for the flat Friedmann-Lemaître-Robertson-Walker (FLRW) solution in radiation domination, the mass within the horizon at the matter-radiation equality $M_{\rm eq}\sim 10^{50}$ g and Eq. (1). This equation is important because it connects the theory and the observation of PBHs. Thus, for example, only a tiny probability $\beta\simeq 2\times 10^{-17}$ is enough to explain all of the dark matter if $M=10^{17}$ g, while $\beta\simeq 2\times 10^{-8}$ is necessary for that if $M=30M_{\odot}$.
@@ 3.1 Overview
The serious study of PBH formation dates back to Carr (1975) Carr (1975). One of its aims is to theoretically predict $\beta(M)$ and, therefore, $f(M)$ and other observables of PBHs for a given cosmological scenario. We can obtain the information of the early Universe from observational data on PBHs only through these studies. Apart from such an observational motivation, through the study of formation we can understand physics in PBH formation and investigate new phenomena and/or new physics in highly nonlinear general relativistic dynamics and high-energy physics. There have been a lot of scenarios proposed for PBH formation. One of the most standard ones is the direct collapse of a large amplitude of primordial perturbations generated by inflation Carr and Lidsey (1993); Carr et al. (1994). See also Ref. Sasaki et al. (2018) for a recent review on this scenario. This can be considered as ‘inevitable’ as the inflationary cosmology has been regarded as an essential part of standard cosmology. Among the other alternative scenarios, domain wall collapse, bubble nucleation, collapse of string networks and phase transitions have attracted more attention than others. Hereafter, we will focus on the formation scenario from fluctuations generated by inflation. Studies on this scenario have developed more than on other scenarios. The key ideas to this scenario is the following: fluctuations generated by inflation, long-wavelength solutions, formation threshold, black hole critical behaviour, dependence on the equation of state (EOS) and statistics on the abundance estimate. We will briefly view these topics below. We will hereafter use the units in which $c=G=1$ over this paper unless they are explicitly given.
@@ 3.2 Fluctuations generated by inflation
The most striking feature of inflation is that it can not only solve the flatness problem and the horizon problem but also provide the mechanism to generate fluctuations through quantum effects. Those fluctuations seed structure formation of different scales in our Universe and provide anisotropies in CMB currently observed with high accuracy.
Figure 1: Schematic figure for the evolution of the scales of fluctuations after generated by inflation. The quantum fluctuations generated by inflation are stretched to the length scales much larger than the Hubble horizon scale $cH^{-1}$ and get classical. In the decelerated phase after inflation, the scale of the fluctuation expands slower than the Hubble horizon scale. The time when the fluctuation scale gets as large as the Hubble scale is called the horizon entry of the perturbation. After the horizon entry, the perturbation can collapse to a black hole leading to PBH formation if its amplitude is sufficiently large.
The evolution of fluctuations in the inflationary cosmology is schematically illustrated in Fig. 1. The quantum fluctuations generated by inflation are stretched to the length scales much larger than the Hubble horizon scale $cH^{-1}$ because the fluctuation scale, which is proportional to $a$, expands faster than the Hubble horizon scale, which is proportional to $t$, due to the accelerated expansion. This process is also considered to remove quantum coherence from the fluctuations. The generated classical fluctuations are further stretched away compared to the Hubble length as long as the inflationary phase continues. After the inflation ends, the expansion of the Universe begins to be decelerated. There is a possibility that the Universe might experience the early matter-dominated phase due to the harmonic oscillation of an inflaton field. In any case, the Universe eventually gets dominated by the radiation field, tightly coupled relativistic particles in an almost complete thermal equilibrium state. This process is called reheating. If we consider the scale of the fluctuation much larger than the Hubble horizon in the decelerated Universe, the scale of the fluctuation, which is proportional to $a$, expands slower than the Hubble horizon scale, which is proportional to $t$. The time when the fluctuation scale gets as large as the Hubble scale is called the horizon entry of the perturbation. Although the inflationary cosmology has been becoming an essential part of the standard cosmology, there has been no standard inflation model until now. There are lots of inflation models, each of which gives the power spectrum $P_{\zeta}(k)$ and the other statistics of the curvature perturbations $\zeta$. See Ref. Sasaki et al. (2018) for details of different inflation models in the context of PBH formation.
@@ 3.3 Large-amplitude long-wavelength solutions
As discussed above, fluctuations are generated and stretched to super-horizon scales by inflation. According to the standard cosmological scenario, the Universe must get radiation-dominated and turn to be decelerated after inflation. Then, the scale of the fluctuations gets as large as the horizon scale, or in other words, the fluctuations re-enter the horizon. Then, if the amplitude of perturbation is nonlinearly large, the fluctuation can collapse to a black hole in the radiation-dominated era. This scenario has been conventionally studied for many years. In some scenarios, there can be an early matter-dominated era before the standard late one. In this case, we also need to think the collapse of the fluctuation in the early matter-dominated era.
Based on the above scenarios, we discuss the fluctuations of super-horizon scales in the decelerated expansion after inflation. This situation is schematically illustrated in Fig. 2.
Figure 2: Schematic figure of long-wavelength solutions. The long-wavelength solutions are obtained under the assumption that the comoving length scale $2\pi/k$ of the perturbation in consideration is much longer than the Hubble horizon length $1/(aH)$, which gives the size of Hubble patches denoted with small blue disks.
As we have already seen, in order to discuss PBH formation in radiation domination, it is necessary to deal with a nonlinearly large amplitude of perturbation of super-horizon scales. To obtain such perturbations, we solve the Einstein equation by a series expansion assuming that the spatial derivative is regarded as much smaller than the time derivative. This scheme is called gradient expansion and the solutions obtained by this scheme are called long-wavelength solutions.
First, we make the standard $3+1$ decomposition of the spacetime. We foliate four-dimensional spacetimes with constant $t$ spacelike hypersurfaces $\Sigma_{t}$. Let $x^{i}$ and $\gamma_{ij}(t,x^{k})$ be spatial coordinates on $\Sigma_{t}$ and the metric tensor on $\Sigma_{t}$, respectively, where $i,j,k,\cdots$ run over 1,2 and 3. Then, the line element in four dimensions can be written by
$ds^{2}=-\alpha^{2}dt^{2}+\gamma_{ij}(dx^{i}+\beta^{i}dt)(dx^{j}+\beta^{j}dt),$ (8)
where $\alpha$ and $\beta^{i}$ are functions of $t$ and $x^{i}$ and called the lapse function and the shift vector, respectively. The functions $\alpha$ and $\beta^{i}$ specify the time slicing, i.e., how the spacetime is foliated with constant $t$ spacelike hypersurfaces, and the coordinate threading, i.e., how the constant $x^{i}$ world lines cross the constant $t$ hypersurfaces, respectively. We can choose the lapse and the shift according to the purposes so that we can fix four degrees of the gauge freedom in the Einstein equation. Using this decomposition, we can reformulate the Einstein equation in the form of the second-order time development, i.e., the Cauchy problem. See e.g. Refs. Wald (1984); Poisson (2009) for more details.
In the next step towards the long-wavelength solutions, we take the following procedures Tomita (1972); Shibata and Sasaki (1999); Lyth et al. (2005); Polnarev and Musco (2007); Harada et al. (2015). In our situation, it is more useful to further decompose $\gamma_{ij}$ in the form
$\gamma_{ij}=e^{2\zeta}a^{2}(t)\tilde{\gamma}_{ij},$ (9)
where $\tilde{\gamma}_{ij}$ and $\zeta$ are both the functions of $t$ and $x^{i}$ so that $\det(\tilde{\gamma}_{ij})=\det(\eta_{ij})$ should hold, where $\eta_{ij}$ is the time-independent metric of the flat 3-space, while $a(t)$ is the scale factor of the reference flat FLRW spacetime. Thus, we have the line element in general spacetimes in four dimensions in the following $3+1$ form:
$ds^{2}=-\alpha^{2}dt^{2}+e^{2\zeta}a^{2}(t)\tilde{\gamma}_{ij}(dx^{i}+\beta^{i}dt)(dx^{j}+\beta^{j}dt).$ (10)
We call this the cosmological $3+1$ conformal decomposition.
Then, we assume that the scale of the perturbation is much larger than the Hubble horizon scale, i.e., that
@@
is much smaller than unity,where $k$ denotes the comoving wave number corresponding to the comoving scale of the perturbation. Applying the gradient expansion in powers of $\epsilon$ for the Einstein equation, we can obtain growing-mode solutions with an assumption that the zeroth-order solutions in powers of $\epsilon$ takes the following form:
$ds^{2}=-dt^{2}+e^{2\zeta_{0}({\bm{x}})}a^{2}(t)\eta_{ij}dx^{i}dx^{j},$ (12)
where $\zeta_{0}({\bm{x}})$ is the zeroth-order part of $\zeta$, which is a function solely of ${\bm{x}}=(x^{1},x^{2},x^{3})$ and identified with primordial curvature perturbation. Note that this assumption is compatible with the comoving slice, the constant-mean-curvature (CMC) slice and the uniform-density slice but not with the conformal Newtonian gauge condition, which is often used in cosmology  [^1] As is well known, in the conformal Newtonian gauge, the density perturbation in the linear order approaches a time-independent function of ${\bm{x}}$ in the super-horizon limit, while it approaches zero in the same limit in the cosmological long-wavelength solutions formulated in Refs. Shibata and Sasaki (1999); Lyth et al. (2005); Polnarev and Musco (2007). We do not exclude the existence of other formulations that are compatible with the conformal Newtonian gauge.. The higher-order terms of the solutions are obtained in terms of the function $\zeta_{0}({\bm{x}})$ if we impose the appropriate gauge conditions. In other words, $\zeta_{0}({\bm{x}})$, which is not suppressed due to the long-wavelength scheme, generates the long-wavelength solutions. For example, the density perturbation $\delta$ is $O(\epsilon^{2})$ and is given by
@@
$ds^{2}=-dt^{2}+e^{2\zeta_{0}(r)}a^{2}(t)[dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})],$ (14)
in the conformally flat coordinates, we can show that this is equivalent to the asymptotically quasihomogeneous solutions developed in Ref. Polnarev and Musco (2007) in the Misner-Sharp formulation, where the comoving slice and the comoving thread are adopted, and the explicit transformation between them is given in Ref. Harada et al. (2015).